Battery life has a strong, negative, linear relationship with temperature. If the least-squares regression line using xx = temperature explains 90% of the variation in battery life, which of the following must be the correlation, rr, between battery life and temperature?

Using the least-squares regression line, y^=−25.5+1.5xy^​=−25.5+1.5x, what is the residual for the data point at (28,19)?
Answer #1

R2 = 0.90

Correlation r = √0.90 = 0.9487

Yhat = -25. 5+1.5X = -25. 5 + 1.5 * 28 = 16.5

Residual = 19 - 16.5= 2.5

Dear student,
I am waiting for your feedback. I have given my 100% to solve your queries. If you satisfied with my answer then please please like this.
Thank You